OpenStudy (anonymous):

Gino designed a garden for his rectangular backyard. The design he drew on paper is 6 inches wide. The width of his backyard is thirty feet. If his drawing is to scale, and his backyard is forty feet long, how many inches long is his drawing? Write a proportion using words. Then, use the proportion to calculate the drawing length. Show your work.

OpenStudy (anonymous):

Giving medal to someone that can help me

jimthompson5910 (jim_thompson5910):

Let x = length of the figure (on paper) \[\Large \frac{\text{distance on paper}}{\text{actual distance}}=\frac{\text{distance on paper}}{\text{actual distance}}\] \[\Large \frac{6 \text{ inches}}{30 \text{ feet}}=\frac{x \text{ inches}}{40 \text{ feet}}\] \[\Large \frac{6}{30}=\frac{x }{40}\] \[\Large 6*40 = 30*x\] keep going to solve for x

OpenStudy (anonymous):

x=8

jimthompson5910 (jim_thompson5910):

x = 8 is correct

jimthompson5910 (jim_thompson5910):

so the length of the figure on paper is 8 inches

OpenStudy (anonymous):

oh i see now thanks

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem